OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (nirmalnema):

y-(-1)=4(x-(-3)) y+1=4x+12 4x-y+11=0

OpenStudy (anonymous):

if you need that in point- slope form its y=4x+11. This is because after you plug in your values into your equation y=mx+b, the new equation is (-1)=4(-3)+b and you solve for b. so: 4(-3)=-12, subtract 12 from both sides and you get 11=b. put the b value in the equation y=mx+b. Your slope (m) is 4, and your b value is 8 (that's what you just solved for).

OpenStudy (anonymous):

my bad, I had a typo you have to add 12 to both sides, but the answer is still the same